OpenStudy (anonymous):

(sq root 6)^8x = 216^x-3 x=-3???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ln (sqrt 6)^(8x) = ln 216^(x–3) Remove exponents and put the in front of the logarithms. 8x * ln sqrt(6) = (x-3) ln 216 Square the square root of six and divide the 8x by two. Factor 216 to 6^3 4x * ln 6 = (x-3) ln 6^3 4x * ln 6 = 3(x-3) ln 6 Divide ln 6 4x = 3x - 9 so: x = -9 make sense?

OpenStudy (anonymous):

I was wrong?!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

FML

OpenStudy (anonymous):

aww.. im sorry..:/

OpenStudy (anonymous):

it's okay thank you for your lovely help.

OpenStudy (anonymous):

np:)